# Forecasting mortality rates with functional signatures

### v1

@@ 3.2.1 A closer look at the French and Australian mortality data
The findings of French and Australian mortality data present a comparative analysis of the HUts model’s performance under differing demographic conditions and historical events influencing mortality data. While the HUts model is able to perform well in stable demographic conditions, its advantages are particularly noticeable when confronting data with potential outliers, making it a robust tool for forecasting when compared with the HU variants and the LC model.
@@ 3.3 Interval forecast
where $\mathbf{1}(\cdot)$ is the indicator function, and $\hat{y}_{t+j|t}^{(\alpha/2)}(x_{i})$ is the $\alpha/2$-quantile from the bootstrapped samples. It quantifies whether the intervals are adequately wide to encompass the actual observations at a specified nominal coverage of $1-\alpha$. 95% and 80% prediction intervals are constructed for all the 12 countries and tabulated in Tables 8 and 9 respectively.
Table 8: Coverage probability of 95% bootstrapped prediction intervals (PI)
Bootstrapped PI (%) Bias-Corrected Bootstrapped PI (%)
Australia 83.86% 88.48%
Belgium 93.38% 95.27%
Bulgaria 88.86% 91.10%
Denmark 76.79% 83.15%
Finland 93.62% 93.12%
France 99.91% 99.85%
Ireland 83.53% 86.66%
Italy 98.94% 98.73%
Japan 96.50% 98.34%
Netherlands 97.11% 97.03%
Norway 84.57% 87.64%
United States 93.17% 95.25%
Table 9: Coverage probability of 80% bootstrapped prediction intervals (PI)
Bootstrapped PI (%) Bias-Corrected Bootstrapped PI (%)
Australia 66.48% 74.25%
Belgium 78.72% 85.46%
Bulgaria 75.91% 78.40%
Denmark 55.97% 65.76%
Finland 82.25% 80.86%
France 97.29% 94.74%
Ireland 64.89% 76.93%
Italy 93.29% 91.24%
Japan 83.81% 95.71%
Netherlands 79.37% 80.04%
Norway 63.56% 72.62%
United States 76.65% 86.42%
@@
Figure 11: Bootstrapped 95% (red dot-dashed line) and 80% (blue dotted line) prediction intervals for one-step-ahead forecast for French mortality (1899-2014) (a) without bias correction (b) with bias correction
@@
Figure 12: Bootstrapped 95% (red dot-dashed line) and 80% (blue dotted line) prediction intervals for one-step ahead forecast for Australian mortality (1974-2014) (a) without bias correction (b) with bias correction
The 95% and 80% bootstrapped prediction intervals are tabulated in Table 8 and 9 respectively. Overall, the bias-correction approach significantly improves the empirical coverage probabilities, bringing them closer to the nominal coverage levels. This adjustment helps to ensure that the prediction intervals are more accurately capturing the true variability in the mortality data. The bias-corrected intervals consistently show coverage probabilities that are closer to the nominal coverage than the bootstrapped intervals without bias-correction, indicating the effectiveness of the correction in improving the reliability of the forecast intervals. For instance, the improvement in empirical coverage is evident in Denmark, where the 95% bias-corrected interval coverage increases from 76.79% to 83.15%, and in Australia, where it rises from 83.86% to 88.48%.
However, bias-correction is not always necessary. In some cases, such as Japan, the empirical coverage probabilities are already close to the nominal levels without bias correction. For Japan, the 95% bootstrapped prediction interval coverage is 96.50%, and the bias-corrected coverage is 98.34%. This suggests that in certain contexts where the initial prediction intervals are already well-calibrated, the additional step of bias correction may not provide significant benefits.
The accuracy of the prediction intervals is closely tied to the accuracy of the point forecasts. Since bias correction centers the intervals around the point forecasts, the quality of these point forecasts directly influences the effectiveness of the interval coverage. Accurate point forecasts lead to well-centered intervals that are more likely to include the actual observed values. Conversely, if the point forecasts are biased or inaccurate, the intervals, even after bias correction, may fail to achieve the desired coverage levels. This interplay between point forecast accuracy and interval coverage will be further illustrated using the mortality data from France and Australia in Figures 11 and 12. In these plots, the black dots represent the observed mortality rates.
The prediction interval plots for France and Australia in reveal significant differences in variability and width, reflecting the nature of each country’s mortality data. France’s data, with notable outliers due to historical events, results in wider, more conservative prediction intervals to capture greater variability. In contrast, Australia’s stable mortality rates lead to narrower intervals; however, less accurate point forecasts, particularly in the 20-40 age range, cause these intervals to miss some observed rates, especially at higher horizons.
@@ 4 Discussion
In mortality modeling and forecasting, the HUts model introduces a novel methodology by taking advantage of the universal nonlinearity of signatures and applying signature regression to decompose mortality curves. Unlike the HU model which relies on the covariance function and is sensitive to outliers, the HUts model learns data dependency through signature coefficients using principal component regression or SVD, effectively reducing the impact of outliers. The principal components obtained thus represent the structure and interactions within the data.
@@
The HUts model has demonstrated an ability to perform competitively with, or even surpass, the other HU model variants across a different scenarios. Notably, the HUts model achieves these results without the need for weighted methods or robust statistics, which are employed by the wHU and HUrob models to improve their performance under specific conditions. This approach of the HUts model delivers robustness and is capable to adapt to diverse datasets. Whether dealing with irregularities as seen in the French mortality data or handling more stable demographic trends like those observed in Australia, the HUts model maintains high accuracy and reliability, making it a highly viable option for actuaries and demographers seeking dependable mortality forecasts without the complexity of additional methodological adjustments.
@@
The HUts model also offers the advantage of easy modification to enhance performance, unlike the HU model’s basis expansion. Therefore, suboptimal results are not indicative of inherent limitations but rather reflect its modular nature, allowing for greater adaptability. For instance, the embedding step is not limited to the basepoint time lead-lag transformation, Morrill et al. (2021) compiled various other augmentations that can be incorporated to improve the model. Additionally, future embeddings may better represent the geometric characteristics of a path. Besides augmentations, Morrill et al. (2021) suggested using a logsignature over the signature to reduce feature vector dimensions, and scaling paths or signatures may also offer improvements.
@@
It is also important to acknowledge the limitations of this study. The prediction intervals constructed under the HUts model face certain limitations, particularly due to the normality assumptions on error terms. Normality assumptions, when not justified, can lead to inaccuracies in the predicted intervals, necessitating the use of bootstrap methods. Bootstrap methods, although useful in addressing non-normality, comes more computational demand. This computational demand can be a significant drawback in practical applications where computational efficiency is crucial. In addition, bootstrap methods are sensitive to outliers which can affect the resampling process, leading to wider prediction intervals. This is evident in the case of French mortality data, which includes periods of war and disease outbreaks. These outliers introduce greater variability into the bootstrap samples, resulting in more conservative and wider intervals. This phenomenon is less pronounced in the mortality data, where the historical data is more stable and follows a general trend without significant outliers.
@@
Future research should focus on refining the HUts model, potentially integrating more advanced smoothing techniques or exploring alternative statistical methods to enhance its performance further. Additionally, the model’s application can be extended to multi-population mortality forecasting as signatures are an excelent dimension reduction tool and incorporating socioeconomic factors as a binary dimension could provide deeper insights into mortality trends. The following practical extensions can be readily implemented:
@@
In addition, the performance of the HUts model, like any other model, depends on the country’s data and forecast horizon, with some models performing relatively better in certain settings. This emphasizes the influence of country-specific mortality dynamics and model sensitivities. As Shang (2012) pointed out, there is no universal mortality model capable of effectively handling data from all countries. This makes the HUts model a candidate for model averaging approaches. Model averaging combines the strengths of various models to improve forecast accuracy by assigning empirical weights to each model based on their performance at different horizons. Given the HUts model’s demonstrated efficacy, especially in long-term forecasts, it can be assigned greater weight in model-averaged forecasts to leverage its strengths (Shang and Booth, 2020). The benefits of model averaging lie in enhancing forecast accuracy through a strategic combination of models. By integrating the HUts model into a model averaging framework, it is possible to achieve more reliable and accurate mortality forecasts, particularly over longer forecast horizons.
Beyond demography, the HUts model extends itself offering a robust framework for addressing a range of challenges in functional data analysis, including functional principal component regression (FPCR), functional regression with functional responses, and forecasting functional time series. The flexibility of the HUts model to adapt to various functional data analysis scenarios is primarily due to its ability to integrate advanced mathematical concepts from functional data analysis, such as smoothing, handling of high-dimensional datasetss. This integration not only enhances the model’s performance in demographic forecasting but also makes it a powerful tool for broader applications in statistical and machine learning fields involving complex, structured data.
5 Conclusion
This paper has introduced the incorporation of signatures into the HU model, resulting in the HUts model. In not only outperforms the HU model but also exhibits enhanced robustness, particularly in the presence of outliers and irregularities in the data. The empirical results from the French and other mortality datasets provide compelling evidence of the HUts model’s efficacy. In the case of France, with the presence of notable historical outliers due to wars and disease outbreaks, the HUts model showed substantial improvement over the HU model. This is evident from the lower MSE and MAE values across various forecast horizons. The enhanced performance can be attributed to the model’s ability to mitigate the influence of outliers, which are more prevalent in the French data. This robustness is crucial for accurate mortality forecasting in regions with volatile historical trends. The HUts model’s superior performance across all age groups, particularly at longer forecast horizons, underscores its utility even in more stable datasets. This consistency across diverse demographic conditions highlights the model’s versatility and reliability. Prediction intervals were constructed with the application of bootstrap procedures. The coverage probabilities of the 80% and 95% bootstrapped prediction intervals with and without bias correction reveal that bias-corrected intervals generally provide better coverage, especially in the absence of outliers. This is particularly pronounced in the French data, where the intervals constructed were more conservative and wide. In contrast, the Australian data, with fewer outliers, shows smaller coverage deviance with its bias corrected bootstrapped prediction interval. From a practical standpoint, these findings have significant implications for actuaries and policymakers. The improved accuracy and robustness of the HUts model can lead to more reliable mortality forecasts, which are essential for various applications, including insurance pricing, pension fund management, and public health planning. The model’s ability to handle data with outliers and provide accurate long-term forecasts makes it a valuable tool for demography.
@@ A.1 Normality assumptions of errors
To obtain prediction intervals using distributional forecasts, all sources of error must satisfy the normality assumptions. If any source of error does not follow a normal distribution, bootstrapping methods should be used instead. The model residual histograms of ages 0, 25, 40 and 70 of the HUts model fitted with French mortality data are plotted with a superimposed normal distribution curve are in Figure 13. It is evident that the residuals do not conform to a normal distribution.
Figure 13: Residual histograms of ages 0, 25, 40 and 70
@@ A.2 Selection of truncation order
A.3 Remaining MSE and MAE plots

### v2

@@ 3.2.1 A closer look at the French and Australian mortality data
The findings of French and Australian mortality data present a comparative analysis of the HUts model’s performance under differing demographic conditions and historical events influencing mortality data. While the HUts model is able to perform well in stable demographic conditions, its advantages are particularly noticeable when confronting data with potential outliers, making it a robust tool for forecasting when compared with the HU model variants.
@@ 3.3 Interval forecast
where $\mathbf{1}(\cdot)$ is the indicator function, and $\hat{y}_{t+j|t}^{(\alpha/2)}(x_{i})$ is the $\alpha/2$-quantile from the bootstrapped samples. It quantifies whether the intervals are adequately wide to encompass the actual observations at a specified nominal coverage of $(1-\alpha)100\%$. 95% prediction intervals of the HUts model and the wHU model are constructed for all the 12 countries and tabulated in Table 2. The minimum observations required to obtain the one-step-ahead forecast errors used for the adjustment, $r$ is set to 30.
Table 2: Empirical coverage probabilities of bootstrapped prediction intervals
HUts wHU
95% Adjusted 95% 95% Adjusted 95%
Australia 92.70% 94.77% 88.30% 89.22%
Belgium 97.36% 96.09% 91.05% 81.13%
Bulgaria 91.89% 85.44% 82.51% 79.47%
Denmark 92.99% 95.84% 78.06% 91.17%
Finland 97.39% 98.80% 94.68% 98.97%
France 100.00% 100.00% 99.95% 100.00%
Ireland 92.81% 82.62% 73.54% 80.54%
Italy 99.84% 99.94% 99.42% 99.86%
Japan 97.08% 97.39% 95.84% 79.40%
Netherlands 99.61% 99.95% 98.15% 99.94%
Norway 94.30% 97.98% 88.26% 98.15%
United States 96.60% 99.11% 92.39% 85.70%
@@
Figure 8: Bootstrapped (a) 95% (b) adjusted 95% prediction intervals of one-step-ahead forecast for French mortality (1899-2014) for the HUts model and the wHU model
@@
Figure 9: Bootstrapped (a) 95% (b) adjusted 95% prediction intervals of one-step-ahead forecast for Australian mortality (1921-2014) for the HUts model and the wHU model
The bootstrapping procedure generally performs well in obtaining prediction intervals that are close to the nominal 95% coverage level. Across various countries and forecasting methods, the unadjusted 95% prediction intervals yield empirical coverage probabilities that align reasonably well with the expected nominal level. These results demonstrate that the bootstrap approach reliably captures the underlying variability in mortality data, ensuring that the prediction intervals accurately achieve the desired 95% coverage.
However, when applying the Hyndman and Shang (2009) adjustment to these prediction intervals, the effectiveness of the adjustment varies across different countries and methods. This adjustment aims to refine the intervals based on one-step-ahead forecast errors, potentially bringing the empirical coverage probabilities closer to the nominal level. Yet, its impact is not uniformly beneficial and depends on the specific characteristics of the mortality data and the forecasting models used.
In the case of Australia, both the prediction intervals of the HUts and wHU models show improvements in coverage probabilities after the adjustment. For the HUts method, the empirical coverage increased from 92.70% to 94.77%, moving closer to the nominal 95% level. Similarly, for the wHU method, the coverage improves from 88.30% to 89.22%. This indicates that the adjustment effectively enhances the reliability of the prediction intervals in these instances. A similar positive effect is observed for Denmark with the wHU model, where the coverage probability significantly increases from 78.06% to 91.17% after adjustment, as well as from 92.99% to 95.84% with the HUts model.
On the other hand, the adjustment does not always lead to better coverage probabilities. For example, in Belgium using the wHU method, the coverage probability decreases from 91.05% to 81.13% after adjustment, moving further away from the nominal level. Similarly, in Bulgaria with the HUts method, the coverage dropped from 91.89% to 85.44% post-adjustment. These instances indicate that the adjustment may not universally improve the prediction intervals possibly due to the nature of the forecast errors of the fitted models.
A notable case is France, where the empirical coverage probabilities reaches 100% for both unadjusted and adjusted intervals with the HUts model. This phenomenon is caused by the prediction intervals absorbing a significant amount of uncertainty from the forecasts, resulting in wider intervals as seen in Figure 8. For the unadjusted prediction intervals, the uncertainty arises from bootstrapping the model errors. Due to the presence of outliers in the French mortality data, these model errors are larger. When the prediction intervals are adjusted, they incorporate additional variability from the in-sample one-step-ahead forecast errors. These forecast errors are also inflated by outliers, resulting in even wider intervals. This further widening makes the predictions more conservative to ensure coverage despite the high variability in the data. The same effect is observed with the wHU model, where the prediction intervals are similarly wide which suggests that outliers can impact the bootstrapping approach, potentially reducing the reliability of the intervals.
In contrast, when examining the prediction intervals for Australia in Figure 9, they are significantly narrower than those for France. Australia’s mortality data is more stable, with fewer outliers, leading to less uncertainty in the forecasts. Consequently, the prediction intervals are narrower. This comparison highlights the point that uncertainty within mortality data leads to wider bootstrapped prediction intervals, as seen in France, while more stable data results in narrower intervals, as observed in Australia.
@@ 4 Discussion
In mortality modeling and forecasting, the HUts model introduces a novel methodology by taking advantage of the universal nonlinearity of signatures and applying signature regression to decompose mortality curves. Unlike the HU model, which relies on the covariance function and is sensitive to outliers, the HUts model learns data dependency through signature coefficients using principal component regression or SVD, effectively reducing the impact of outliers. The principal components obtained thus represent the structure and interactions within the data.
@@
The HUts model performs competitively and often surpasses other HU model variants in various scenarios. Notably, the HUts model achieves these results without the need for weighted methods or robust statistics, which are employed by the wHU and HUrob models to improve their performance under specific conditions. This approach of the HUts model delivers robustness and is capable of adapting to diverse datasets. Whether dealing with irregularities as seen in the French mortality data or handling more stable demographic trends like those observed in Australia, the HUts model maintains high accuracy and reliability, making it a highly viable option for actuaries and demographers seeking dependable mortality forecasts without the complexity of additional methodological adjustments.
@@
The HUts model also offers the advantage of easy modification to enhance performance, unlike the HU model’s basis expansion. Therefore, suboptimal results are not indicative of inherent limitations but rather reflect its modular nature, allowing for greater adaptability. For instance, the embedding step is not limited to the basepoint time lead-lag transformation, Morrill et al. (2021) compiled various other augmentations that can be incorporated to improve the model and better represent the geometric characteristics of a path. Besides augmentations, applying the log-signature (Morrill et al., 2021) or a randomized signature (Cuchiero et al., 2020; Compagnoni et al., 2023) (see A.7 in the Appendix) over the signature can also be considered to reduce the feature vector dimensions. Morrill et al. (2021) also suggests that the scaling of paths or signatures may offer improvements.
@@
It is also important to acknowledge the limitations of this study. The prediction intervals constructed under the HUts model face certain limitations, particularly due to the normality assumptions on error terms. Normality assumptions, when not justified, can lead to inaccuracies in the predicted intervals, necessitating the use of bootstrap methods. Bootstrap methods, although useful in addressing non-normality, come with more computational demand. This computational demand can be a significant drawback in practical applications where computational efficiency is crucial. In addition, bootstrap methods are sensitive to outliers which can affect the resampling process, leading to wider prediction intervals. This is evident in the case of French mortality data, which includes periods of war and disease outbreaks. These outliers introduce greater variability into the bootstrap samples, resulting in more conservative and wider intervals. This phenomenon is less pronounced in the mortality data, where the historical data is more stable and follows a general trend without significant outliers. Steps such as adapting the robust bootstrap procedure from Beyaztas and Shang (2022) can be taken to reduce the influence of outliers by assigning weights to model residuals.
In addition, the performance of the HUts model, like any other model, depends on the country’s data and forecast horizon, with some models performing relatively better in certain settings. This emphasizes the influence of country-specific mortality dynamics and model sensitivities. As Shang (2012) pointed out, there is no universal mortality model capable of effectively handling data from all countries. This makes the HUts model a candidate for model averaging approaches. Model averaging combines the strengths of various models to improve forecast accuracy by assigning empirical weights to each model based on their performance at different horizons. Given the HUts model’s demonstrated efficacy, especially in long-term forecasts, it can be assigned greater weight in model-averaged forecasts to leverage its strengths (Shang and Booth, 2020; Chang and Shi, 2023). The benefits of model averaging lie in enhancing forecast accuracy through a strategic combination of models. By integrating the HUts model into a model averaging framework, it is possible to achieve more reliable and accurate mortality forecasts, particularly over longer forecast horizons.
Beyond demography, the HUts model extends itself by offering a robust framework to address a range of challenges in functional data analysis, including functional principal component regression (FPCR), functional regression with functional responses, and forecasting functional time series. The flexibility of the HUts model to adapt to various functional data analysis scenarios is primarily due to its ability to integrate advanced mathematical concepts from functional data analysis, such as smoothing and handling of high-dimensional datasets. This integration not only enhances the model’s performance in demographic forecasting but also makes it a powerful tool for broader applications in statistical and machine learning fields involving complex, structured data.
5 Conclusion
This paper has introduced the incorporation of signatures into the HU model, resulting in the HUts model. It is able to outperform the HU model variants while exhibiting enhanced robustness, particularly in the presence of outliers and irregularities in the data. The empirical results from the French and other mortality datasets provide compelling evidence of the HUts model’s efficacy. In the case of France, with the presence of notable historical outliers due to wars and disease outbreaks, the HUts model showed substantial improvement over the HU model variants. This is evident from the lower MSE values across various forecast horizons. The enhanced performance can be attributed to the model’s ability to mitigate the influence of outliers, which are more prevalent in the French data. This robustness is crucial for accurate mortality forecasting in regions with volatile historical trends. The HUts model’s superior performance across all age groups, particularly at longer forecast horizons, displays the model’s versatility across diverse demographic conditions. Prediction intervals were constructed with the application of bootstrap procedures. The coverage probabilities of the 95% bootstrapped prediction intervals generally are close to the intended nominal coverage, while the adjustment is able to improve said coverage for certain cases. The presence of outliers can impact the construction of the intervals, which is particularly pronounced in the case of French mortality, where the intervals constructed were more conservative and wide. In contrast, Australian mortality, with fewer outliers, shows smaller coverage deviance with its bootstrapped prediction interval. From a practical standpoint, these findings have significant implications for actuaries and policymakers. The improved accuracy and robustness of the HUts model can lead to more reliable mortality forecasts, which are essential for various applications, including insurance pricing, pension fund management, and public health planning. The model’s ability to handle data with outliers and provide accurate long-term forecasts makes it a valuable tool for demography.
@@ 5 Conclusion
Future research should focus on refining the HUts model, potentially integrating more advanced smoothing techniques or exploring alternative statistical methods to enhance its performance further. Additionally, the model’s application can be extended to multi-population mortality forecasting as signatures are an excellent dimension reduction tool, and incorporating socioeconomic factors as a binary dimension could provide deeper insights into mortality trends. The following practical extensions can be readily implemented:
@@ A.1 Normality assumptions of errors
To obtain prediction intervals using distributional forecasts, all sources of error must satisfy the normality assumptions. If any source of error does not follow a normal distribution, bootstrapping methods should be used instead. While it is unrealistic to expect perfect normality, we prefer that most of the residuals resemble a normal distribution; otherwise, bootstrapped methods are preferred. We have plotted the residual histograms and QQ plots of ages 0, 10, 25, 50, 75, and 100 for all four models fitted with French (1899-2014) mortality data . Only the residuals of the HU model, seen in Figure 11, seem to resemble somewhat of a normal distribution. The residuals from the three remaining models, seen in Figures 10, 12, and 13, deviate significantly from a normal distribution and would benefit from the bootstrapping procedure when constructing prediction intervals.
(a) Histogram
(b) QQ-plot
Figure 10: HUts model residuals
(a) Histogram
(b) QQ-plot
Figure 11: HU model residuals
(a) Histogram
(b) QQ-plot
Figure 12: HUrob model residuals
(a) Histogram
(b) QQ-plot
Figure 13: wHU model residuals
@@ A.2 Selection of truncation order
A.3 Data availability and codes
The codes are available upon request.
A.4 Mean squared error and mean absolute error tables
HUts HU HUrob wHU
Australia 0.00982 0.01027 0.01696 0.00952
Belgium 0.02183 0.02144 0.02419 0.02047
Bulgaria 0.01942 0.02131 0.02158 0.01968
Denmark 0.03636 0.03810 0.04098 0.03590
Finland 0.03856 0.04271 0.04435 0.03705
France 0.00398 0.00642 0.00710 0.00399
Ireland 0.05062 0.05174 0.05735 0.05161
Italy 0.00711 0.00920 0.01600 0.00680
Japan 0.00674 0.00747 0.00938 0.00687
Netherlands 0.01242 0.01344 0.01569 0.01143
Norway 0.04397 0.04303 0.04800 0.04264
United States 0.00133 0.00207 0.00535 0.00129
Mean 0.02101 0.02227 0.02558 0.02061
HUts HU HUrob wHU
Australia 0.01835 0.02245 0.02890 0.01642
Belgium 0.02627 0.02863 0.03211 0.02620
Bulgaria 0.04595 0.05212 0.04574 0.05478
Denmark 0.04891 0.05484 0.05764 0.05007
Finland 0.04698 0.05988 0.05328 0.04538
France 0.00882 0.02290 0.01489 0.00978
Ireland 0.07035 0.07198 0.07707 0.07785
Italy 0.01625 0.03153 0.02982 0.01546
Japan 0.01045 0.01288 0.01361 0.01020
Netherlands 0.01785 0.02171 0.02374 0.01859
Norway 0.05056 0.05309 0.05856 0.05053
United States 0.00746 0.00900 0.01252 0.00896
Mean 0.03068 0.03675 0.03732 0.03202
HUts HU HUrob wHU
Australia 0.0418 0.0511 0.0509 0.0351
Belgium 0.0364 0.0380 0.0450 0.0404
Bulgaria 0.0976 0.0931 0.0811 0.1280
Denmark 0.0764 0.0825 0.0927 0.0825
Finland 0.0653 0.0894 0.0773 0.0617
France 0.0201 0.0526 0.0278 0.0242
Ireland 0.1111 0.1297 0.1177 0.1413
Italy 0.0403 0.0603 0.0594 0.0433
Japan 0.0348 0.0399 0.0289 0.0362
Netherlands 0.0326 0.0391 0.0406 0.0377
Norway 0.0695 0.0775 0.0802 0.0718
United States 0.0226 0.0196 0.0342 0.0305
Mean 0.0540 0.0644 0.0613 0.0611
The mean absolute error expresses the average difference between predicted and observed values and is computed using:
$MAE(h)=\frac{1}{pq}\sum^{q}_{t=1}\sum^{p}_{i=1}\lvert y_{t}(x_{i})-\hat{y}_{t|t-h}(x_{i})\rvert.$
HUts HU HUrob wHU
Australia 0.06111 0.06614 0.09498 0.06036
Belgium 0.08404 0.09092 0.10116 0.08193
Bulgaria 0.09106 0.10052 0.09727 0.09138
Denmark 0.11169 0.12029 0.12909 0.11127
Finland 0.11279 0.12608 0.13552 0.10976
France 0.04246 0.06100 0.06336 0.04256
Ireland 0.12512 0.13419 0.14838 0.12719
Italy 0.05330 0.06842 0.08936 0.05210
Japan 0.05061 0.05701 0.06955 0.04975
Netherlands 0.06802 0.07839 0.08692 0.06657
Norway 0.11718 0.11890 0.13299 0.11479
United States 0.02648 0.03459 0.05380 0.02628
Mean 0.07866 0.08804 0.10020 0.07783
HUts HU HUrob wHU
Australia 0.09131 0.11006 0.12925 0.08682
Belgium 0.10163 0.11696 0.12482 0.10307
Bulgaria 0.15511 0.17190 0.15103 0.15636
Denmark 0.14989 0.16413 0.17112 0.15113
Finland 0.14040 0.17654 0.15844 0.13800
France 0.07102 0.12234 0.09527 0.07427
Ireland 0.16031 0.18212 0.18869 0.17752
Italy 0.09528 0.14200 0.13301 0.08831
Japan 0.07659 0.08745 0.08784 0.07277
Netherlands 0.09511 0.11317 0.11668 0.09702
Norway 0.13801 0.15506 0.16166 0.13688
United States 0.06428 0.07333 0.08508 0.06843
Mean 0.11158 0.13459 0.13357 0.11255
HUts HU HUrob wHU
Australia 0.1483 0.1705 0.1749 0.1383
Belgium 0.1323 0.1421 0.1558 0.1434
Bulgaria 0.2490 0.2389 0.2093 0.2341
Denmark 0.2097 0.2182 0.2352 0.2151
Finland 0.1804 0.2334 0.2105 0.1754
France 0.1150 0.1905 0.1360 0.1204
Ireland 0.2164 0.2663 0.2510 0.2641
Italy 0.1623 0.2056 0.2019 0.1550
Japan 0.1410 0.1536 0.1289 0.1435
Netherlands 0.1448 0.1601 0.1637 0.1564
Norway 0.1809 0.2089 0.2087 0.1837
United States 0.1121 0.1122 0.1380 0.1250
Mean 0.1660 0.1917 0.1845 0.1712
A.5 Remaining MSE and MAE plots